OpenStudy (dontknowdontcare):

3. There are two fruit trees located at (3,0) and (−3, 0) in the backyard plan. Maurice wants to use these two fruit trees as the focal points for an elliptical flowerbed. Johanna wants to use these two fruit trees as the focal points for some hyperbolic flowerbeds. Create the location of two vertices on the y-axis. Show your work creating the equations for both the horizontal ellipse and the horizontal hyperbola. Include the graph of both equations and the focal points on the same coordinate plane.

OpenStudy (dontknowdontcare):

@Cas_fangirl_14

OpenStudy (cas_fangirl_14):

I'm stuck, sorry. :(

OpenStudy (dontknowdontcare):

It's okay I'm stuck too lol

OpenStudy (cas_fangirl_14):

@ShadowLegendX @heyyyyitsholly can you help?

OpenStudy (dontknowdontcare):

@mrm

OpenStudy (mrm):

Do you know the equation for an ellipse?

OpenStudy (mrm):

The equations for an ellipse and a hyperbola completely depends on the coordinates of the foci, these can be looked up in your book, or found with a quick google search.

OpenStudy (dontknowdontcare):

c^2=a^2-b^2 is the equation of an ellipse

OpenStudy (dontknowdontcare):

and then for a hyperbola it is c^2=a^2+b^2

OpenStudy (mrm):

Use equations that are in Cartesian form, meaning they are in terms of x and y.

OpenStudy (adrimit):

I think he is referring to the axis lengths